OpenStudy (anonymous):

what is the solution of the system of equations 2x -5y = 11 and -2 + 3y = -9? This is a multiple choice question: (1). (-3,-1) (2). (-1,3) (3). (3,-1) (4). (3,1)

OpenStudy (anonymous):

Substitute each ordered pair into either of the equations and solve, the one that gives you consistent results, is the correct answer: Pair 1: 2(-3)-5(-1)=11 True? Pair 2: 2(-1)-5(3)=11 True? Pair 3: 2(3)-5(-1)=11 True? Pair 4: 2(3)-5(1)=11 True? The one where you answered true is correct. To be 100% certain, do that for both equations, the correct result must be consistent for both equations.

OpenStudy (anonymous):

Oh yeah! i never thought of that.